OpenStudy (anonymous):

The distance between Milltown and summerville is 3.5 inches on a map. If the map scale is 1 inch = 20 miles, what is the actual distance? Plz help i will Medal and fan

OpenStudy (anonymous):

@xo.minnie.xox plz help

OpenStudy (anonymous):

70

OpenStudy (anonymous):

Can you explain to me @becca1

OpenStudy (anonymous):

hi @FrazCraz

OpenStudy (anonymous):

hi @Rubyjadeluv

OpenStudy (anonymous):

can someone help me on my question

OpenStudy (anonymous):

1 inch = 20 miles 2 inch = 40 miles 3 inch = 60 miles .5 inch = 10 miles

OpenStudy (anonymous):

does that help

OpenStudy (anonymous):

yes thanks :)

OpenStudy (anonymous):

can i get a medal

OpenStudy (anonymous):

no cuz i dont understand @becca1

OpenStudy (anonymous):

the actuall distance is 70? how did you find that im trying to say lol

OpenStudy (anonymous):

1 inch is 20 miles if you do 3 times 20 you get 60 you still got a half half of 20 is ten then do 60 plus 10 and you get 70 miles

OpenStudy (anonymous):

ok thanks i faned and medaled

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

@becca1 what do u mean i stil have a half?

OpenStudy (anonymous):

i need help:(

OpenStudy (anonymous):

me to

OpenStudy (anonymous):

3*20=60 1/2*20=10 60+10=70

OpenStudy (anonymous):

now thats more like it :) ty lol @becca1